4(n+3)/4=n+3
multi step algebra

Answer by unlockmath(1688)

Hello,
With this 4(n+3)/4=n+3
The 4s cancel out to give us:
n+3=n+3
From this we can see any number can be replaced for "n" therefore the answers are infinite.
Make sense?
RJ
www.math-unlock.com